# A Survey of AI Agent Protocols

### v1

@@ Abstract
The rapid development of large language models (LLMs) has led to the widespread deployment of LLM agents across diverse industries, including customer service, content generation, data analysis, and even healthcare. However, as more LLM agents are deployed, a major issue has emerged: there is no standard way for these agents to communicate with external tools or data sources. This lack of standardized protocols makes it difficult for agents to work together or scale effectively, and it limits their ability to tackle complex, real-world tasks. A unified communication protocol for LLM agents could change this. It would allow agents and tools to interact more smoothly, encourage collaboration, and triggering the formation of collective intelligence. In this paper, we provide a systematic overview of existing communication protocols for LLM agents. We classify them into four main categories and make an analysis to help users and developers select the most suitable protocols for specific applications. Additionally, we conduct a comparative performance analysis of these protocols across key dimensions such as security, scalability, and latency. Finally, we explore future challenges, such as how protocols can adapt and survive in fast-evolving environments, and what qualities future protocols might need to support the next generation of LLM agent ecosystems. We expect this work to serve as a practical reference for both researchers and engineers seeking to design, evaluate, or integrate robust communication infrastructures for intelligent agents.
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), LLM agents [^1] For presentation brevity, in this paper, the multi-modal LLM concept (Caffagni et al., 2024) is merged into the LLM concept. are increasingly being deployed across various industries, including automated customer service, content creation, data analysis, and medical assistance (OpenAI et al., 2024; Gottweis et al., 2025; Lin et al., 2025; Yang et al., 2025a; Zhang et al., 2024a; Guo et al., 2024; Zhou et al., 2024), transforming our daily work and life. To fully exploit the potential of agents, many architectures have emerged to facilitate communication between agents and external entities. These entities include resources not directly controlled by agents, such as various data sources and tools, as well as other online agents.
@@ 2.1 Definition and Characteristics of LLM agents
  • Memory Systems: LLM agents implement both short-term and long-term memory components to maintain context across interactions and store relevant information for future use (Zhang et al., 2024b; Yang et al., 2025b). This dual memory system allows agents to maintain conversation continuity while building knowledge over time.
@@ 3.2.1 General-Purpose Protocols
A2A (Agent2Agent Protocol) Complex Problem Solving of Agents Intra-Enterprise Agent Collaboration RPC, OAuth Landing
@@
Agent2Agent (A2A) Protocol is a kind of agent collaboration protocol proposed by Google, designed to enable seamless agent collaboration regardless of underlying frameworks and vendor implementations. It simplifies the integration of agents within enterprise environments and provides core functions required to buildecure, enterprise-grade agent ecosystem. These functions include capability discovery, user experience negotiation, task and state management, and secure collaboration. Therefore, A2A is specifically designed to support complex intra-organizational agent collaboration. The key principles of ANP are:
@@
  • Enterprise Readiness: The protocol is designed with built-in considerations for authentication, authorization, security, privacy, traceability, and observability. Agents are treated as enterprise-grade applications, ensuring robustness and security in production environments.
@@ 4.8 Evaluation over Protocol Evolution: Case Studies
  • Contextual Fit: Select the appropriate protocol based on the specific application scenario. For example, MCP is ideal for scenarios requiring integration with external tools and data sources; ANP is more suitable for cross-domain communication and collaboration among agents on the Internet; A2A offers more comprehensive support for intra-organizational agent collaboration.
@@ 5.1 MCP: Single Agent Invokes All Tools
5.2 A2A: Complex Collaboration Within an Enterprise
Agent-to-Agent (A2A) protocol enables direct communication between agents within an enterprise’s boundary. As depicted in Figure 4 (upper right), the A2A implementation distributes intelligence across multiple specialized agents:
@@ 5.2 A2A: Complex Collaboration Within an Enterprise
A2A protocol demonstrates a more flexible and realistic enterprise architecture where services can establish direct connections when needed. For instance, the A2A Flight Agent can obtain weather information directly from the Weather Agent without going through the Travel Planner. This reduces unnecessary communication overhead and allows for more complex collaboration patterns.
@@ 5.3 ANP: Cross-Domain Agent Protocol
Agent Network Protocol (ANP) extends collaboration beyond enterprise boundaries through standardized cross-domain interactions. Figure 4 (lower left) shows how ANP enables negotiations between agents in different organizational domains:
@@
ANP addresses the challenges of cross-organizational collaboration by formalizing the negotiation process. Unlike A2A, which assumes trust within an enterprise, ANP establishes clear protocols for cross-domain requests and responses. This makes it particularly suitable for scenarios involving multiple organizations with distinct business interests and security boundaries.
@@ 5.4 Agora: Natural Language to Protocol Generation
Through the case study, it becomes evident that each protocol has specific conditions and dependencies for successful application. 1) MCP requires a centralized agent (e.g., travel assistant) and stable external servers (flight, hotel, weather) with well-defined interfaces, making it ideal for well-defined tasks but limited in dynamic environments. 2) A2A, on the other hand, is suitable for enterprise settings where multiple specialized agents (e.g., flight, hotel, weather) communicate directly. Its success depends on a stable internal infrastructure and efficient data exchange among agents, without needing centralized coordination. 3) ANP extends collaboration across organizational boundaries, requiring each entity (e.g., airline, hotel, weather service) to follow its clear protocols for data exchange, making it ideal for cross-organizational collaboration where negotiations and security protocols are essential. 4) Finally, Agora focuses on user experience by translating natural language into structured protocols, requiring a system capable of understanding user inputs and mapping them to appropriate service agents. Each protocol’s applicability is shaped by factors such as organizational structure, trust relationships, and the need for efficient user interaction, highlighting the importance of selecting the right protocol based on the specific requirements of the use case.

### v2

@@ Abstract
The rapid development of large language models (LLMs) has led to the widespread deployment of LLM agents across diverse industries, including customer service, content generation, data analysis, and even healthcare. However, as more LLM agents are deployed, a major issue has emerged: there is no standard way for these agents to communicate with external tools or data sources. This lack of standardized protocols makes it difficult for agents to work together or scale effectively, and it limits their ability to tackle complex, real-world tasks. A unified communication protocol for LLM agents could change this. It would allow agents and tools to interact more smoothly, encourage collaboration, and triggering the formation of collective intelligence. In this paper, we provide the first comprehensive analysis of existing agent protocols, proposing a systematic two-dimensional classification that differentiates context-oriented versus inter-agent protocols and general-purpose versus domain-specific protocols. Additionally, we conduct a comparative performance analysis of these protocols across key dimensions such as security, scalability, and latency. Finally, we explore the future landscape of agent protocols by identifying critical research directions and characteristics necessary for next-generation protocols. These characteristics include adaptability, privacy preservation, and group-based interaction, as well as trends toward layered architectures and collective intelligence infrastructures. We expect this work to serve as a practical reference for both researchers and engineers seeking to design, evaluate, or integrate robust communication infrastructures for intelligent agents.
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), LLM agents [^1] For presentation brevity, in this paper, the multi-modal LLM concept (Caffagni et al., 2024) is merged into the LLM concept. are increasingly being deployed across various industries, including automated customer service, content creation, data analysis, and medical assistance (OpenAI et al., 2024; Gottweis et al., 2025; Yang et al., 2025a; Guo et al., 2024; Zhou et al., 2024), transforming our daily work and life. To fully exploit the potential of agents, many architectures have emerged to facilitate communication between agents and external entities. These entities include resources not directly controlled by agents, such as various data sources and tools, as well as other online agents.
@@ 2.1 Definition and Characteristics of LLM agents
  • Memory Systems: LLM agents implement both short-term and long-term memory components to maintain context across interactions and store relevant information for future use (Zhang et al., 2024; Yang et al., 2025b). This dual memory system allows agents to maintain conversation continuity while building knowledge over time.
@@ 3.2.1 General-Purpose Protocols
A2A (Agent2Agent Protocol) Complex Problem Solving of Agents Inter Agent Collaboration RPC, OAuth Landing
@@
Agent2Agent (A2A) Protocol is a kind of agent collaboration protocol proposed by Google, designed to enable seamless agent collaboration regardless of underlying frameworks and vendor implementations. It simplifies the integration of agents within different environments and provides core functions required to buildecure, like enterprise-grade agent ecosystem. These functions include capability discovery, user experience negotiation, task and state management, and secure collaboration. Therefore, A2A is specifically designed to support complex inter agent collaboration. The key principles of ANP are:
@@
  • Enterprise Readiness: The protocol is designed with built-in considerations for authentication, authorization, security, privacy, traceability, and observability. Agents can be treated as enterprise-grade applications, ensuring robustness and security in production environments.
@@ 4.8 Evaluation over Protocol Evolution: Case Studies
  • Contextual Fit: Select the appropriate protocol based on the specific application scenario. For example, MCP is ideal for scenarios requiring integration with external tools and data sources; ANP is more suitable for cross-domain communication and collaboration among agents on the Internet; A2A offers more comprehensive support for inter agent collaboration.
@@ 5.1 MCP: Single Agent Invokes All Tools
5.2 A2A: Complex Collaboration Inter-agents Within an Enterprise
Agent-to-Agent (A2A) protocol enables direct communication between different agents for complex tasks. As depicted in Figure 4 (upper right), the A2A implementation distributes intelligence across multiple specialized agents:
@@ 5.2 A2A: Complex Collaboration Inter-agents Within an Enterprise
A2A protocol demonstrates a more flexible and realistic architecture where agents can establish direct connections when needed. For instance, the A2A Flight Agent can obtain weather information directly from the Weather Agent without going through the Travel Planner. This reduces unnecessary communication overhead and allows for more complex collaboration patterns across any type of organization or multi-agent system.
@@ 5.3 ANP: Cross-Domain Agent Protocol
Agent Network Protocol (ANP) extends collaboration through standardized cross-domain interactions. Figure 4 (lower left) shows how ANP enables negotiations between agents in different organizational domains:
@@
ANP addresses the challenges of collaboration between independent agents by formalizing the protocol-based interaction process. While A2A focuses on message-based delegation, ANP establishes clear protocols for structured requests and responses between agents. This makes it particularly suitable for scenarios involving agents with distinct capabilities, well-defined interfaces, and potentially different security boundaries, regardless of whether they exist within the same system or across multiple systems.
@@ 5.4 Agora: Natural Language to Protocol Generation
Through the case study, it becomes evident that each protocol has specific conditions and dependencies for successful application. 1) MCP employs a centralized agent (e.g., travel assistant) that sequentially invokes tools with clear interfaces to accomplish tasks. This approach works efficiently for well-defined workflows but may require central agent modification to adapt to new scenarios. 2) A2A enables collaboration through message/data exchange between specialized agents (e.g., flight, hotel, weather). Each agent autonomously handles its assigned task and communicates results back to a coordinating agent, allowing for flexible communication patterns while maintaining overall coordination. 3) ANP utilizes structured protocol-based interactions where the primary agent retains processing logic but delegates specific execution steps through well-defined API-like interfaces. This standardized approach works effectively regardless of whether agents exist within the same system or across different domains. 4) Finally, Agora focuses on translating natural language into appropriate structured protocols, serving as an intermediary layer that maps user intentions to the specific protocols required by different agents. Each protocol’s applicability depends on factors such as the desired level of agent autonomy, communication flexibility, interface standardization, and the complexity of the tasks being performed.